OpenStudy (anonymous):

A school band sold 30 raffle tickets, each labeled with a number from 1 to 30. One winning ticket will be drawn. What is the probability that the number of the winning ticket will be a multiple of 4 or the number 19? Express your answer in simplest form.

OpenStudy (alivejeremy):

need help

OpenStudy (anonymous):

what i can guess

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

4/15

OpenStudy (alivejeremy):

There are 7 numbers from 1 to 30 that are multiples of 4: 4, 8, 12, 16, 20, 24, 28 So there are 8 ways that a number can be 19 or a multiple of 4. So the probability of that is 8/30, which is 4/15 in simplest form.

OpenStudy (alivejeremy):

how many multiples of 4 are there between 1 and 30 ? 4, 8, 12, 16, 20, 24, 28 there are 7 multiples of 4 and 1 number 19 there are 8 ways to win P=8/30=4/15

OpenStudy (anonymous):

do u know that @siblings is my sister

OpenStudy (anonymous):

6th

OpenStudy (alivejeremy):

no lol

OpenStudy (alivejeremy):

0 possible answers 7 tickets = multiples of 4 (4,8,12,16,20,24,28) 1 ticket = 19 therefore: P(multiple of 4 or 19) = 7+1 / 30 = 8/30 = 4/15 ... think that's the simplest fraction you'll get, you can work that out in decimals or % tho if u like?

OpenStudy (anonymous):

after i am done with this i have 1 more lesson semester

OpenStudy (anonymous):

a medal 4 u

OpenStudy (alivejeremy):

bk to u

OpenStudy (anonymous):

i will make a new Q

OpenStudy (alivejeremy):

lol

OpenStudy (anonymous):

k